### v1

# Perturbation Analysis of Singular Values in Concatenated Matrices

Abstract
Concatenating matrices is a common technique for uncovering shared structures in data through singular value decomposition (SVD) and low-rank approximations. However, a fundamental question arises: how does the singular value spectrum of the concatenated matrix relate to the spectra of its individual components? In this work, we develop a perturbation framework that extends classical results such as Weyl’s inequality to concatenated matrices. We establish analytical bounds that quantify the stability of singular values under small perturbations in the submatrices. Our results show that if the matrices being concatenated are close in norm, the dominant singular values of the concatenated matrix remain stable, enabling controlled trade-offs between accuracy and compression. These insights provide a theoretical foundation for improved matrix clustering and compression strategies, with applications in numerical linear algebra, signal processing, and data-driven modeling.
1   Introduction
In this work, we focus on the analysis of concatenated matrices. A concatenated matrix is defined as a matrix formed by horizontally (or vertically) stacking a collection of submatrices. In particular, given $N$ matrices $\{A_{i}\}_{i=1}^{N}$, each with the same number of rows, their horizontal concatenation is given by
$A=[A_{1}\;A_{2}\;\cdots\;A_{N}].$
This construction allows us to jointly analyze multiple matrices by extracting a common basis, which is especially useful for revealing shared structures in different data representations.
The singular value decomposition (SVD) is a cornerstone of numerical linear algebra, serving as a powerful tool for dimensionality reduction, principal component analysis, and data compression [6, 3]. In many modern applications, data is naturally represented by multiple matrices – each representing different aspects of an underlying phenomenon. A common strategy to reveal the intrinsic structure of such data is to concatenate these matrices and extract a shared basis through SVD.
An important aspect of our study is how the number of submatrices influences the compression performance. When compressing $N$ matrices, one can either compress each matrix individually or concatenate a subset (or all) of them and then compress the resulting block matrix. The key idea is that by compressing a concatenated matrix, one can identify and exploit a common basis among the submatrices, leading to a higher compression rate. On the other hand, compressing the matrices separately may result in a lower compression rate since the redundancy and shared information among the submatrices are not leveraged as effectively.
However, an important question arises: how do small perturbations in the individual matrices influence the singular values of the concatenated matrix and, consequently, the stability of the shared basis?
Classical perturbation theory provides valuable insights into this question. Foundational results such as Weyl’s theorem [9], along with further developments in eigenvector and singular value perturbation [2, 8, 7], offer estimates on the sensitivity of spectral properties to perturbations. In this work, we extend these classical results to the setting of concatenated matrices by deriving analytical bounds that explicitly quantify how perturbations in individual blocks affect the singular value spectrum of the concatenated matrix. Our analysis confirms the intuition that if the submatrices are close in norm, the dominant singular values of the concatenated matrix remain stable – a property that can be used to control the trade-off between compression accuracy and the number of matrices being grouped.
Evidently, these results have immediate practical implications. Namely, modern algorithms for randomized matrix decompositions [5] and clustering [4, 1] can benefit from our perturbation estimates by providing a guiding framework for concatenating matrices. Such concatenation is essential when designing algorithms that must balance computational efficiency with the preservation of intrinsic data structure, especially in high-dimensional and large-scale settings.
2   Preliminaries
We begin with a classical result in perturbation theory which gives a uniform bound on the change of eigenvalues of symmetric matrices when they are perturbed.
Theorem 1 (Weyl’s Inequality [9]).
Let $A,B\in\mathbb{R}^{n\times n}$ be symmetric matrices and define the perturbation $E=B-A$. Then for every eigenvalue (ordered in any fixed manner) we have
$|\lambda_{i}(B)-\lambda_{i}(A)|\leq\|E\|_{2},$
where $\|\cdot\|_{2}$ denotes the spectral (operator) norm.
Weyl’s inequality is a cornerstone in matrix perturbation theory. It assures us that if the perturbation $E$ is small in the spectral norm, then the eigenvalues of the perturbed matrix $B$ will remain close to those of the original matrix $A$. This property becomes crucial when we later examine how singular values are affected by perturbations in concatenated matrices.
Next, we review the relationship between the singular values of a matrix and the eigenvalues of its Gram matrix.
Proposition 1.
For any matrix $A\in\mathbb{R}^{n\times m}$ with rank $r$, the singular values of $A$ are given by
$\sigma_{i}(A)=\sqrt{\lambda_{i}(A^{T}A)},\quad 1\leq i\leq r.$
Proof.
Let
$A=U\Sigma V^{T},$
be the singular value decomposition (SVD) of $A$, where $U\in\mathbb{R}^{n\times n}$ and $V\in\mathbb{R}^{m\times m}$ are orthogonal matrices, and $\Sigma\in\mathbb{R}^{n\times m}$ is a diagonal matrix containing the singular values $\sigma_{i}(A)$. Then,
$A^{T}A=V\Sigma^{T}U^{T}U\Sigma V^{T}=V(\Sigma^{T}\Sigma)V^{T}.$
Since $\Sigma^{T}\Sigma$ is a diagonal matrix with diagonal entries $\sigma_{1}^{2}(A),\sigma_{2}^{2}(A),\dots,\sigma_{r}^{2}(A)$ (with the remaining entries being zero), the eigenvalues of $A^{T}A$ are exactly these squares. Thus,
$\sigma_{i}(A)=\sqrt{\lambda_{i}(A^{T}A)}\quad\text{for }1\leq i\leq r.$
An analogous argument holds for $AA^{T}$, thus confirming the result. ∎
We now focus on a specific matrix construction that is central to our work. Consider a matrix formed by the concatenation of the same copy of matrix $A$. The structure of the constructed matrix induces a certain regularity that we can exploit.
Proposition 2.
Let $A\in\mathbb{R}^{m\times n}$ be a matrix with $\operatorname{rank}(A)=r$. Define the concatenated matrix
$M=\begin{pmatrix}A&A&\dots&A\end{pmatrix}\in\mathbb{R}^{m\times kn},$
which consists of $k$ copies of $A$. Define
$S_{1}=M^{T}M\quad\text{and}\quad S_{2}=MM^{T}.$
Then:
  • The nonzero eigenvalues of both $S_{1}$ and $S_{2}$ are given by
    $\lambda_{i}(S_{1})=\lambda_{i}(S_{2})=k\,\lambda_{i}(A^{T}A),\quad 1\leq i\leq r.$
  • The remaining eigenvalues are zero:
    $\lambda_{i}(S_{1})=0\quad\text{for }r<i\leq kn,$
    $\lambda_{i}(S_{2})=0\quad\text{for }r<i\leq m.$
Proof.
Observe that
$S_{2}=MM^{T}=\begin{pmatrix}A&A&\dots&A\end{pmatrix}\begin{pmatrix}A^{T}\\ A^{T}\\ \vdots\\ A^{T}\end{pmatrix}=k\,AA^{T}.$
If $x_{i}$ is an eigenvector of $A^{T}A$ corresponding to the eigenvalue $\lambda_{i}(A^{T}A)$, then by the properties of the SVD, the nonzero eigenvalues of $AA^{T}$ are the same as those of $A^{T}A$. Consequently, for $i=1,\dots,r$,
$S_{2}\,x_{i}=k\,AA^{T}\,x_{i}=k\,\lambda_{i}(A^{T}A)\,x_{i}.$
Thus, $\lambda_{i}(S_{2})=k\,\lambda_{i}(A^{T}A)$ for $1\leq i\leq r$, with the remaining eigenvalues being zero. Since $M^{T}M$ and $MM^{T}$ share the same nonzero eigenvalues, the claim for $S_{1}$ follows. ∎
Finally, we present a useful bound for block matrices, which will be used when dealing with perturbations structured in a block format.
Proposition 3 (Bounding the Norm of a Block Matrix).
Let
$E=\begin{pmatrix}E_{11}&E_{12}&\dots&E_{1k}\\ E_{21}&E_{22}&\dots&E_{2k}\\ \vdots&\vdots&\ddots&\vdots\\ E_{k1}&E_{k2}&\dots&E_{kk}\end{pmatrix}$
be a block matrix. Then,
$\|E\|_{2}\leq\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{k}\|E_{ij}\|_{2}^{2}},$
where $\|\cdot\|_{2}$ denotes the spectral norm.
Proof.
For any nonzero vector $x$, partition it conformably as
$x=\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{k}\end{pmatrix}.$
Then,
$Ex=\begin{pmatrix}\sum_{j=1}^{k}E_{1j}x_{j}\\[2.84526pt] \sum_{j=1}^{k}E_{2j}x_{j}\\[2.84526pt] \vdots\\[2.84526pt] \sum_{j=1}^{k}E_{kj}x_{j}\end{pmatrix}.$
Define $y_{i}=\sum_{j=1}^{k}E_{ij}x_{j}$ for $i=1,\dots,k$. By the triangle inequality and the submultiplicative property of the spectral norm, we have
$\|y_{i}\|_{2}\leq\sum_{j=1}^{k}\|E_{ij}x_{j}\|_{2}\leq\sum_{j=1}^{k}\|E_{ij}\|_{2}\|x_{j}\|_{2}.$
Using the Cauchy–Schwarz inequality,
$\|y_{i}\|_{2}^{2}\leq\left(\sum_{j=1}^{k}\|E_{ij}\|_{2}^{2}\right)\left(\sum_{j=1}^{k}\|x_{j}\|_{2}^{2}\right).$
Since $\|x\|_{2}^{2}=\sum_{j=1}^{k}\|x_{j}\|_{2}^{2}$, summing over $i$ yields
$\|Ex\|_{2}^{2}=\sum_{i=1}^{k}\|y_{i}\|_{2}^{2}\leq\left(\sum_{i=1}^{k}\sum_{j=1}^{k}\|E_{ij}\|_{2}^{2}\right)\|x\|_{2}^{2}.$
Taking the square root and then maximizing over all $x\neq 0$ gives
$\|E\|_{2}\leq\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{k}\|E_{ij}\|_{2}^{2}}.$
3   Bound Estimation
In this section, we establish an upper bound on the perturbation error for a concatenated matrix. We begin by introducing the necessary definitions and notation.
Let $\{A_{i}\}_{i=1}^{k}$ be a collection of matrices with $A_{i}\in\mathbb{R}^{m\times n}$ for each $i\in[1,k]$. Define the original concatenated matrix as
$M=[A_{1},A_{2},\dots,A_{k}],$
and the perturbed concatenated matrix as
$\widetilde{M}=[\widetilde{A}_{1},\widetilde{A}_{2},\dots,\widetilde{A}_{k}].$
The perturbation is represented by
$E=\widetilde{M}-M=[E_{1},E_{2},\dots,E_{k}],$
where $E_{i}=\widetilde{A}_{i}-A_{i}$ for each $i\in[1,k]$.
We now present key lemmas that characterize the structure of the perturbation.
Lemma 1.
Using the above definitions, we have the identity
$\widetilde{M}^{T}\widetilde{M}=M^{T}M+D,$
where the perturbation term $D$ satisfies the bound
$\|D\|_{2}\leq\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{k}\big{(}\|A_{i}\|_{2}\|E_{j}\|_{2}+\|E_{i}\|_{2}\|A_{j}\|_{2}+\|E_{i}\|_{2}\|E_{j}\|_{2}\big{)}^{2}}.$
Proof.
Since each perturbed matrix can be written as $\widetilde{A}_{i}=A_{i}+E_{i}$, it follows that
$\widetilde{M}^{T}\widetilde{M}=\begin{bmatrix}(A_{1}+E_{1})^{T}\\ \vdots\\ (A_{k}+E_{k})^{T}\end{bmatrix}[(A_{1}+E_{1}),\dots,(A_{k}+E_{k})].$
Expanding the product,
$\widetilde{M}^{T}\widetilde{M}=\begin{bmatrix}(A_{1}+E_{1})^{T}(A_{1}+E_{1})&\dots&(A_{1}+E_{1})^{T}(A_{k}+E_{k})\\ \vdots&\ddots&\vdots\\ (A_{k}+E_{k})^{T}(A_{1}+E_{1})&\dots&(A_{k}+E_{k})^{T}(A_{k}+E_{k})\end{bmatrix}.$
Each block takes the form
$(A_{i}+E_{i})^{T}(A_{j}+E_{j})=A_{i}^{T}A_{j}+A_{i}^{T}E_{j}+E_{i}^{T}A_{j}+E_{i}^{T}E_{j}.$
Thus, the unperturbed term is
$M^{T}M=\begin{bmatrix}A_{1}^{T}A_{1}&\dots&A_{1}^{T}A_{k}\\ \vdots&\ddots&\vdots\\ A_{k}^{T}A_{1}&\dots&A_{k}^{T}A_{k}\end{bmatrix},$
and the perturbation matrix $D$ is given by
$D=\begin{bmatrix}A_{1}^{T}E_{1}+E_{1}^{T}A_{1}+E_{1}^{T}E_{1}&\dots&A_{1}^{T}E_{k}+E_{1}^{T}A_{k}+E_{1}^{T}E_{k}\\ \vdots&\ddots&\vdots\\ A_{k}^{T}E_{1}+E_{k}^{T}A_{1}+E_{k}^{T}E_{1}&\dots&A_{k}^{T}E_{k}+E_{k}^{T}A_{k}+E_{k}^{T}E_{k}\end{bmatrix}.$
Defining the individual block components as $D_{ij}=A_{i}^{T}E_{j}+E_{i}^{T}A_{j}+E_{i}^{T}E_{j}$, we obtain
$D=\begin{bmatrix}D_{11}&\dots&D_{1k}\\ \vdots&\ddots&\vdots\\ D_{k1}&\dots&D_{kk}\end{bmatrix}.$
Applying the submultiplicative property of the norm,
$\|D_{ij}\|_{2}\leq\|A_{i}^{T}E_{j}\|_{2}+\|E_{i}^{T}A_{j}\|_{2}+\|E_{i}^{T}E_{j}\|_{2}\leq\|A_{i}\|_{2}\|E_{j}\|_{2}+\|E_{i}\|_{2}\|A_{j}\|_{2}+\|E_{i}\|_{2}\|E_{j}\|_{2}.$
From Proposition 3, it follows that
$\|D\|_{2}\leq\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{k}\|D_{ij}\|_{2}^{2}},$
which gives the bound
$\|D\|_{2}\leq\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{k}\big{(}\|A_{i}\|_{2}\|E_{j}\|_{2}+\|E_{i}\|_{2}\|A_{j}\|_{2}+\|E_{i}\|_{2}\|E_{j}\|_{2}\big{)}^{2}}.$
Lemma 2.
Using the above definitions, the following holds:
$\widetilde{M}\widetilde{M}^{T}=MM^{T}+D$
where the perturbation term $D$ satisfies the bound
$\|D\|_{2}\leq\sum_{i=1}^{k}\left(2\|A_{i}\|_{2}\|E_{i}\|_{2}+\|E_{i}\|_{2}^{2}\right).$
Proof.
We begin by expressing $\widetilde{M}\widetilde{M}^{T}$ using the concatenated form of $\widetilde{M}$:
$\widetilde{M}\widetilde{M}^{T}=[\widetilde{A}_{1},\dots,\widetilde{A}_{k}]\begin{bmatrix}\widetilde{A}_{1}^{T}\\ \vdots\\ \widetilde{A}_{k}^{T}\end{bmatrix}.$
This gives us the sum
$\widetilde{M}\widetilde{M}^{T}=\sum_{i=1}^{k}\widetilde{A}_{i}\widetilde{A}_{i}^{T}.$
Next, recall that $\widetilde{A}_{i}=A_{i}+E_{i}$ for all $i\in[1,k]$, so we can expand $\widetilde{A}_{i}\widetilde{A}_{i}^{T}$ as follows:
$\widetilde{A}_{i}\widetilde{A}_{i}^{T}=(A_{i}+E_{i})(A_{i}+E_{i})^{T}=A_{i}A_{i}^{T}+A_{i}E_{i}^{T}+E_{i}A_{i}^{T}+E_{i}E_{i}^{T}.$
Thus, we can write the sum as
$\widetilde{M}\widetilde{M}^{T}=\sum_{i=1}^{k}\left(A_{i}A_{i}^{T}+A_{i}E_{i}^{T}+E_{i}A_{i}^{T}+E_{i}E_{i}^{T}\right).$
We now compare this with the original matrix $MM^{T}=\sum_{i=1}^{k}A_{i}A_{i}^{T}$, yielding:
$\widetilde{M}\widetilde{M}^{T}=MM^{T}+\sum_{i=1}^{k}\left(A_{i}E_{i}^{T}+E_{i}A_{i}^{T}+E_{i}E_{i}^{T}\right).$
Thus, the perturbation term $D$ is given by
$D=\sum_{i=1}^{k}\left(A_{i}E_{i}^{T}+E_{i}A_{i}^{T}+E_{i}E_{i}^{T}\right).$
To bound $\|D\|_{2}$, we apply the triangle inequality for the spectral norm:
$\|D\|_{2}\leq\sum_{i=1}^{k}\|A_{i}E_{i}^{T}+E_{i}A_{i}^{T}+E_{i}E_{i}^{T}\|_{2}.$
We now apply the subadditivity property of the spectral norm:
$\|A_{i}E_{i}^{T}+E_{i}A_{i}^{T}+E_{i}E_{i}^{T}\|_{2}\leq\|A_{i}E_{i}^{T}\|_{2}+\|E_{i}A_{i}^{T}\|_{2}+\|E_{i}E_{i}^{T}\|_{2}.$
Next, we estimate each term:
$\|A_{i}E_{i}^{T}\|_{2}\leq\|A_{i}\|_{2}\|E_{i}\|_{2},\quad\|E_{i}A_{i}^{T}\|_{2}\leq\|E_{i}\|_{2}\|A_{i}\|_{2},\quad\|E_{i}E_{i}^{T}\|_{2}\leq\|E_{i}\|_{2}^{2}.$
Therefore, we obtain the following bound for $\|D\|_{2}$:
$\|D\|_{2}\leq\sum_{i=1}^{k}\left(2\|A_{i}\|_{2}\|E_{i}\|_{2}+\|E_{i}\|_{2}^{2}\right).$
Through Lemmas 1 and 2, we derived explicit bounds for the perturbations in the matrices $M^{T}M$ and $MM^{T}$, which serve as the key components in understanding the stability of these matrices under small changes.
The results demonstrate that the perturbation in $MM^{T}$ is less significant compared to $M^{T}M$, as it involves only diagonal blocks in its perturbation structure. This distinction provides a tighter bound for $MM^{T}$, offering stronger stability guarantees.
4   Main Result
Theorem 2 (Singular Value Perturbation of a Concatenated Matrix).
Let $\{A_{i}\}_{i=1}^{k}$ be a collection of matrices with $A_{i}\in\mathbb{R}^{m\times n}$ for all $i\in[1,k]$. Define the original concatenated matrix by
$M=[A_{1},\dots,A_{k}]\in\mathbb{R}^{m\times kn},$
and denote its rank by $r=\operatorname{rank}(M)$. Let $\widetilde{M}=[\widetilde{A}_{1},\dots,\widetilde{A}_{k}]$ be the perturbed version of $M$, and define the perturbation matrix
$E=\widetilde{M}-M=[E_{1},\dots,E_{k}],\quad\text{with }E_{i}=\widetilde{A}_{i}-A_{i}\text{ for each }i\in[1,k].$
Then, the following perturbation bounds for the singular values hold:
  • For $i=1,\dots,r$ (corresponding to the nonzero singular values of $M$),
    $|\sigma_{i}(\widetilde{M})-\sigma_{i}(M)|\leq\frac{1}{\sigma_{i}(M)}\sum_{j=1}^{k}\Bigl{(}2\|A_{j}\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)}.$
  • For $i=r+1,\dots,\min(m,kn)$ (corresponding to the zero singular values of $M$),
    $\sigma_{i}(\widetilde{M})\leq\sqrt{\sum_{j=1}^{k}\Bigl{(}2\|A_{j}\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)}}.$
Proof.
We begin by applying Weyl’s inequality to two related symmetric matrices, namely $M^{T}M$ and $MM^{T}$, to obtain bounds on the perturbations of their eigenvalues, which are related to the singular values of $M$ and $\widetilde{M}$.
Step 1: Analysis via $M^{T}M$.
By Lemma 1, we can express
$\widetilde{M}^{T}\widetilde{M}=M^{T}M+D_{1},$
with the perturbation term $D_{1}$ satisfying
$\|D_{1}\|_{2}\leq\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{k}\Bigl{(}\|A_{i}\|_{2}\|E_{j}\|_{2}+\|E_{i}\|_{2}\|A_{j}\|_{2}+\|E_{i}\|_{2}\|E_{j}\|_{2}\Bigr{)}^{2}}.$
Applying Weyl’s inequality to the eigenvalues of $M^{T}M$ and $\widetilde{M}^{T}\widetilde{M}$, we obtain
$|\lambda_{i}(\widetilde{M}^{T}\widetilde{M})-\lambda_{i}(M^{T}M)|\leq\|D_{1}\|_{2}.$
Since, by Proposition 1, the eigenvalues of $M^{T}M$ and $\widetilde{M}^{T}\widetilde{M}$ are the squares of the singular values, i.e.,
$\lambda_{i}(\widetilde{M}^{T}\widetilde{M})=\sigma_{i}^{2}(\widetilde{M})\quad\text{and}\quad\lambda_{i}(M^{T}M)=\sigma_{i}^{2}(M),$
this yields
$|\sigma_{i}^{2}(\widetilde{M})-\sigma_{i}^{2}(M)|\leq\|D_{1}\|_{2}.$
Step 2: Analysis via $MM^{T}$.
Similarly, by Lemma 2 we have
$\widetilde{M}\widetilde{M}^{T}=MM^{T}+D_{2},$
where
$\|D_{2}\|_{2}\leq\sum_{i=1}^{k}\Bigl{(}2\|A_{i}\|_{2}\|E_{i}\|_{2}+\|E_{i}\|_{2}^{2}\Bigr{)}.$
Weyl’s inequality then implies that
$|\lambda_{i}(\widetilde{M}\widetilde{M}^{T})-\lambda_{i}(MM^{T})|\leq\|D_{2}\|_{2}.$
Again, using Proposition 1 we note that
$\lambda_{i}(\widetilde{M}\widetilde{M}^{T})=\sigma_{i}^{2}(\widetilde{M})\quad\text{and}\quad\lambda_{i}(MM^{T})=\sigma_{i}^{2}(M),$
so that
$|\sigma_{i}^{2}(\widetilde{M})-\sigma_{i}^{2}(M)|\leq\|D_{2}\|_{2}.$
Since the bound provided by Lemma 2 (and hence $\|D_{2}\|_{2}$) is tighter than that from Lemma 1, we will use $\|D_{2}\|_{2}$ in the remainder of the proof.
Deriving the Final Bound.
For indices $i=1,\dots,r$, where $\sigma_{i}(M)>0$, note that
$\sigma_{i}^{2}(\widetilde{M})-\sigma_{i}^{2}(M)=\bigl{(}\sigma_{i}(\widetilde{M})-\sigma_{i}(M)\bigr{)}\bigl{(}\sigma_{i}(\widetilde{M})+\sigma_{i}(M)\bigr{)}.$
Since singular values are nonnegative and $\sigma_{i}(\widetilde{M})+\sigma_{i}(M)>0$, it follows that
$|\sigma_{i}(\widetilde{M})-\sigma_{i}(M)|\leq\frac{\|D_{2}\|_{2}}{\sigma_{i}(\widetilde{M})+\sigma_{i}(M)}.$
Moreover, because $\sigma_{i}(\widetilde{M})+\sigma_{i}(M)>\sigma_{i}(M)$, we obtain
$|\sigma_{i}(\widetilde{M})-\sigma_{i}(M)|\leq\frac{\|D_{2}\|_{2}}{\sigma_{i}(M)}\leq\frac{1}{\sigma_{i}(M)}\sum_{j=1}^{k}\Bigl{(}2\|A_{j}\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)}.$
For indices $i=r+1,\dots,\min(m,kn)$, we have $\sigma_{i}(M)=0$. In this case,
$\sigma_{i}^{2}(\widetilde{M})\leq\|D_{2}\|_{2},$
which implies
$\sigma_{i}(\widetilde{M})\leq\sqrt{\|D_{2}\|_{2}}\leq\sqrt{\sum_{j=1}^{k}\Bigl{(}2\|A_{j}\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)}}.$
We now explore two important corollaries that illustrate applications of Theorem 2 in specific settings.
4.1   Perturbation Around the Centroid (Clusterization Case)
In many clustering and centroid analysis applications, one encounters the case where the original matrix is formed by repeating a single matrix, and perturbations occur only in some of the copies. More precisely, let
$A\in\mathbb{R}^{m\times n}$
with rank $r=\operatorname{rank}(A)$, and consider perturbations $\{E_{i}\}_{i=2}^{k}$ with $E_{i}\in\mathbb{R}^{m\times n}$ for $i=2,\dots,k$. Define the original concatenated matrix by
$M=[A,A,\dots,A]\in\mathbb{R}^{m\times kn},$
and let the perturbed matrix be given by
$\widetilde{M}=[A,A+E_{2},A+E_{3},\dots,A+E_{k}]\in\mathbb{R}^{m\times kn}.$
Note that here $E_{1}\equiv 0$.
Corollary 1 (Perturbation Around the Centroid).
Then, for $i=1,\dots,r$,
$|\sigma_{i}(\widetilde{M})-\sqrt{k}\,\sigma_{i}(A)|\leq\frac{1}{\sqrt{k}\,\sigma_{i}(A)}\sum_{j=2}^{k}\Bigl{(}2\|A\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)},$
and for $i=r+1,\dots,\min(m,kn)$,
$\sigma_{i}(\widetilde{M})\leq\sqrt{\sum_{j=2}^{k}\Bigl{(}2\|A\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)}}.$
Proof.
The result follows directly from Theorem 2. In this setting, since $A_{j}=A$ for all $j$ and $E_{1}\equiv 0$, the bound in Theorem 2 becomes
$|\sigma_{i}(\widetilde{M})-\sigma_{i}(M)|\leq\frac{1}{\sigma_{i}(M)}\sum_{j=2}^{k}\Bigl{(}2\|A\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)}.$
Furthermore, from Propositions 1 and 2 it follows that
$\sigma_{i}^{2}(M)=\lambda_{i}(M^{T}M)=k\,\lambda_{i}(A^{T}A)=k\,\sigma_{i}^{2}(A),$
so that
$\sigma_{i}(M)=\sqrt{k}\,\sigma_{i}(A).$
Substituting this into the previous inequality yields the stated result. ∎
4.2   Continuity of the Singular Values under Small Perturbations
An immediate consequence of Corollary 1 is a continuity result. Namely, if each perturbation is small, then the singular values of the perturbed matrix converge to the singular values of the original (scaled) matrix.
Corollary 2.
Suppose that $\|E_{j}\|_{2}<\epsilon$ for all $j$. Then,
$\lim_{\epsilon\to 0}\sigma_{i}(\widetilde{M})=\sqrt{k}\,\sigma_{i}(A)\quad\text{for }1\leq i\leq r,$
and
$\lim_{\epsilon\to 0}\sigma_{i}(\widetilde{M})=0\quad\text{for }r<i\leq\min(m,kn).$
Proof.
From Corollary 1, for $i=1,\dots,r$ we have
$|\sigma_{i}(\widetilde{M})-\sqrt{k}\,\sigma_{i}(A)|$ $\leq\frac{1}{\sqrt{k}\,\sigma_{i}(A)}\sum_{j=2}^{k}\Bigl{(}2\|A\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)}$
$\leq\frac{1}{\sqrt{k}\,\sigma_{i}(A)}\sum_{j=2}^{k}\Bigl{(}2\|A\|_{2}\epsilon+\epsilon^{2}\Bigr{)}$
$=\frac{(k-1)\epsilon}{\sqrt{k}\,\sigma_{i}(A)}\Bigl{(}2\|A\|_{2}+\epsilon\Bigr{)}.$
Thus, as $\epsilon\to 0$, it follows that
$\lim_{\epsilon\to 0}|\sigma_{i}(\widetilde{M})-\sqrt{k}\,\sigma_{i}(A)|=0,$
so that
$\lim_{\epsilon\to 0}\sigma_{i}(\widetilde{M})=\sqrt{k}\,\sigma_{i}(A).$
For $i=r+1,\dots,\min(m,kn)$, Corollary 1 implies
$\sigma_{i}(\widetilde{M})$ $\leq\sqrt{\sum_{j=2}^{k}\Bigl{(}2\|A\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\Bigr{)}}$
$\leq\sqrt{\sum_{j=2}^{k}\Bigl{(}2\|A\|_{2}\epsilon+\epsilon^{2}\Bigr{)}}$
$=\sqrt{(k-1)\epsilon\Bigl{(}2\|A\|_{2}+\epsilon\Bigr{)}}.$
Hence, as $\epsilon\to 0$, we obtain
$\lim_{\epsilon\to 0}\sigma_{i}(\widetilde{M})=0.$
Acknowledgments
The authors confirm that there is no conflict of interest and acknowledges financial support by the Simons Foundation grant (1290607, A.T.) and the project 0125U000299 of the National Academy of Sciences of Ukraine. We also express our gratitude to the Armed Forces of Ukraine for their protection, which has made this research possible.